2. The total cost equation of a firm is given by the equation

  TC=5000+2000Q-10Q2+0.5Q3

where TC is total cost and Q is the level of output.

a)     What is the firm’s total fixed cost?

b)     What is the equation for the firm’s total variable cost (TVC)?

c)     What is the equation for the firm’s average total cost (ATC)?

d)    What is the equation for the firm’s marginal cost (MC)?


1
Expert's answer
2021-06-21T12:02:41-0400


Given,

 TC=5000+2000Q-10Q2+0.5Q3

a)

Total Fixed Cost will be

"5000"


b)

"TVC=2000-10Q^2+0.5Q^3"


c)

"ATC=\frac{TC}{Q}"

"ATC=\frac{5000+2000Q-10Q^2+0.5Q^3}{Q}"

"=\frac{5000}{Q}+2000-10Q+0.5Q^2"


d)

"MC=\frac{\delta TC}{\delta Q}"

"MC=\frac{TC=5000+2000Q-10Q^2+0.5Q^3}{\delta Q}"

"=2000-20Q+1.5Q^2"